I am trying to solve the linear system of equations by graphing. the problem reads x/3+y/5=7/15 and the second part states x/4+3/8y=1/8. I know I must find a common denominator and put the problems into y=mx+b form. I am just having a hard time understanding them.
Answer by mananth(16946)

x/3+y/5=7/15
multiply by 15
5x+3y= 7
3y=-5x+7
y=(-5/3)x+7/3.............................(2)

x/4+3/8y=1/8
multiply by 8
2x+3y=1
3y=-2x+1
y=(-2/3)x+1/3...........................(2)

Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 2.33 ,( 0 , 2.33 )
when y = 0 x= 7 ,( 1.4 , 0 )
Equation 2
when x= 0 y= 0.33 ( 0 , 0.33 )
when y = 0 x= 0.5 ( 0 , 0.5 )
you may take simpler values to graph easily.


(2,-1)